OpenStudy (anonymous):

Helpful steps please, Solve the equation in the interval [0°,360°]. sin 2(theta) = -sin(theta)

OpenStudy (welshfella):

use the identity sin 2(theta) = 2 sin(theta) cos(theta)

OpenStudy (anonymous):

I'm still stuck I'm not sure how to progress using that identity.

OpenStudy (welshfella):

i'll write x instead of theta right as I am a bit lazy! 2 sin x cos x = - sin x add sin x to both sides 2 sin x cos x + sin x = 0 now you can factor the left hand side , then its easy

OpenStudy (anonymous):

would it then be 0,120,180,240 ?

OpenStudy (welshfella):

sin x(2 cos x + 1 ) = 0 so sin x = 0 or 2cos x + 1 = 0 sin x = 0 or cos x = -1/2

OpenStudy (welshfella):

for sin x = 0 x = 0, 180,360 for cos x = -1/2 x is in the 2nd and 3rd quadrants

OpenStudy (anonymous):

ok so cos=-1/2 x would be 60,120,240,300 but for sin x=0 they only have 0 and 180 not 360 for the set of solutions

OpenStudy (anonymous):

the interval is [0 degrees, 360 degrees) *

OpenStudy (welshfella):

no cos is negative only in the 2nd and 3rd quadrants its 120 and 240 only

OpenStudy (welshfella):

no 360 is a solution the squared brackets means you include 360

OpenStudy (welshfella):

- oh it looks like ] to me

OpenStudy (anonymous):

my mistake I naturally did the " ]" instead of " )" by habit

OpenStudy (welshfella):

but if its ) then yes - no 360

OpenStudy (welshfella):

so 0,120,180,240 is the correct solution

OpenStudy (anonymous):

ah thanks!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

hace you seen this memory aid